OpenStudy (anonymous):

find the sum of the finite geometric series by using the formular for Sn 1-1/3+1/9-1/27+1/81-1/243+1/729

OpenStudy (anonymous):

divide the sequence as follows... [1+1/9+1/81+1/729]-[1/3+1/27+1/243] thes r 2 diff gps .... solve for their sum as individuals nd subtrat it

OpenStudy (anonymous):

to find the finite

OpenStudy (anonymous):

yes there r three terms in each ... => n=3 r in each case is 1/9 a in case 1 is = 1 nd case 2 = 1/3 use theformula Sn = a[1-(r^n)]/(1-r)

OpenStudy (anonymous):

I got it thank u soo much

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

You could also use (1-r^n)/(1-r) with a ratio of -1/3 (a is 1) which will give u the same result.